[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jun 15, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jun 15 17:52:25 UTC 2017


HSU Shift Log for Thu Jun 15 2017 13:52:17 GMT-0400 (EDT)

Shift: H

Forecaster: J Cangialosi

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: GOES-16 CONUS VIS red band has stripes of missing data.  
It was particularly bad around 14-15 UTC.

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170615/5a4d0ea6/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list